Each month, an American household generates an average of 30 pounds of newspaper
for garbage or recycling. Assume the standard deviation is 5 pounds. If a household is
selected at random, find the probability of its generating
a. Between 25 and 35 pounds per month
b. More than 40 pounds per month
Assume the variable is approximately normally distributed

Answer by mathmate(429)

mean = 30 lbs per month
standard deviation = 5 lbs
A random selection will give
a. 25-35 lbs per month
z1 = (25-30)/5 = -1
z2 = (35-30)/5 = +1
P(z1<=x<=z2)
= P(x = 0.8413 - 0.1587
= 0.6827
======
b. more than 40 lbs per month
z2 = (40-30)/5 = 2
P(x>=z2)
= 1-P(x<=z2)
= 1-0.9772
= 0.02275
=======